OpenStudy (anonymous):

pleaseee helpp!! A person is watching a boat from the top of a lighthouse. The boat is approaching the lighthouse directly. When first noticed the angle of depression to the boat is 18°33'. When the boat stops, the angle of depression is 51°33'. The lighthouse is 200 feet tall. How far did the boat travel from when it was first noticed until it stopped? Round your answer to the hundredths place.

OpenStudy (anonymous):

i tried it and i got tan 51 33'= 200/base1= 158.802 tan(18 33')=200/base2 =596.008 distance travelled= 596.008-158.802=438m

OpenStudy (anonymous):

but my teacher told me there was a rounding error

OpenStudy (wolf1728):

My God, that teacher is a stickler. Yes, you do have an error in your calculations. You stated that 596.008-158.802=438m Actually 596.008-158.802= 437.206 You were in error by almost 80 centimeters! Anyway, if you are to round it to the nearest hundredth then your answer should be 437.21 meters.